OpenStudy (agl202):

Medal Fan for best response!!! Write an equation of a line perpendicular to y = x + 4 in slope-intercept form that passes through the point (-2, 6). y= x - 4 y= x + 6 y= -4x - 2 y= -4x +

OpenStudy (yolo_king):

y = -x

OpenStudy (vivek3461):

Line in slope intercept form is y = mx + c, where 'm' is slope and c is ' Y intercept. condition for perpendicular lines are m1.m2 = -1 where m1, m2 are the slope of perpendicular line. find the slope using above condition and substitute the given point to find the y intercept.

OpenStudy (agl202):

@TrojanPoem How abt u?

OpenStudy (trojanpoem):

Something is wrong with me i got - before the y

OpenStudy (agl202):

So, is it a, b, c, or d???

OpenStudy (yolo_king):

Its C.

OpenStudy (trojanpoem):

Retype the choices

OpenStudy (agl202):

y= 1/4x - 4 y= 1/4x + 6 y= -4x - 2 y= -4x + 1/4

OpenStudy (vivek3461):

Choices seem to be all wrong

OpenStudy (trojanpoem):

Yeah

OpenStudy (yolo_king):

Are these the choices? A) \[y=\frac{ 1 }{ 4 }x-4\] B)\[y=\frac{ 1 }{ 4 }x+6\] C) y= -4x - 2 D) y= -4x + 1/4

OpenStudy (agl202):

Ya...:)

OpenStudy (vivek3461):

None of the slope of the lines given in the choices satisfies the perpendicularity condition

OpenStudy (agl202):

Soooooooo.......

OpenStudy (trojanpoem):

Recheck the question.

OpenStudy (bohotness):

wait what are you looking for?

OpenStudy (trojanpoem):

He is looking for the equation of the line which is prep to the one given.